# Non-asymptotic Estimates for Markov Transition Matrices with Rigorous Error Bounds

### v2

@@ 3.3. Spectral properties of $\bm{P}_{2}$ and $\tilde{\bm{P}}_{2}$
{choose $\mathbb{P}\left\{\text{choose }v\right\}=p(u,v),$ (3.24)
@@
$-\frac{1}{4}\left(\bm{T}_{1}-\bm{S}_{1}^{\mathrm{T}}\right)\left(\bm{T}_{1}^{\mathrm{T}}-\bm{S}_{1}\right)=\frac{1}{2}\left(\bm{D}_{\mu}^{1/2}\frac{\bm{P}-\bm{\mathbf{I}}}{2}\bm{D}_{\mu}^{-1/2}+\bm{D}_{\mu}^{-1/2}\frac{\bm{P}^{\mathrm{T}}-\bm{\mathbf{I}}}{2}\bm{D}_{\mu}^{1/2}\right).$ (3.32)
@@
Similar to the proof for $\bm{P}_{2}$, we decompose $\tilde{\bm{P}}_{2}$ as the product of two transition matrices which represent step 1 and step 2 introduced in Subsection 3.1.2, respectively. To be specific, let $\tilde{\bm{S}}$ denote the transition matrix of step 1 in which a given state $\tilde{\bm{u}}\sim(u,v)$ will jump to $u$ or $v$ with the equal probability $1/2$, i.e. for any $\tilde{\bm{u}}\in\tilde{\Omega}^{2}$,$u\in\Omega$,
@@
{choose $\mathbb{P}\left\{\text{choose }v\right\}=p(u,v),$ (3.45)
@@ 4.1. Tail bounds of MLE and SCE
We will first derive tail bounds for the error of MLE and SCE using concentration inequalities for Markov chains. Let us state the complete version of [10, Theorem 2.4], which is a Bernstein-type inequality for Markov sums that only requires a nonzero IP gap of the transition matrix.
@@
In the above theorem, we use the symbol $\bm{A}$ for an arbitrary Markov transition kernel, so that it will not be confused with the particular transition matrix $\bm{P}$ in our context. Since $\eta_{p}(\bm{A})\geq\eta_{s}(\bm{A})\geq\eta_{a}(\bm{A})$ and $\eta_{p}(\bm{A})\geq\eta_{ps}(\bm{A})/2$, this concentration inequality also holds if the transition admits a nonzero absolute spectral gap, a nonzero pseudo spectral gap,or a nonzero symmetric spectral gap. We then use Theorem 4.1 to derive non-asymptotic tail bounds for MLE.
@@
As for SCE, we choose to derive non-asymptotic tail bounds directly in the matrix form using matrix concentration inequalities. Recently, [19] established Hoeffding and Bernstein-type matrix concentration inequalities (in $l_{2}$ operator norm) for sums of Markov dependent random matrices, under the assumption that the transition matrix admits a nonzero absolute spectral gap. We present a simplified version of [19, Theorem 2.2] as follows.
@@ 4.2. Dimension-free expectation bound
Whether the error depends on the dimension shall rely on the choice of the error metric. The reason we can obtain dimension free expectation bound is that we measure the error using the Frobenius norm (i.e. $L_{2,u})$, which is a natural choice for matrices. If one instead uses the TV norm (i.e. $L_{1,u}$) to measure the error, a dependence on the dimension typically arises. This is why in [23], the error bound depends on the quantity $\|\bm{D}_{\mu}\bm{P}\|_{1/2}=(\sum_{u,v\in\Omega}(\mu(u)p(u,v))^{1/2})^{2}$, which may depend on the dimension $d$ and can be bounded by $d^{2}$.
@@
The estimate in (4.24) can be view as a matrix version of [5, Theorem 1.2].In fact, one can apply [5, Theorem 1.2] to $\{\bm{u}_{k}\}_{k\geq 1}$ to obtain that
@@
In this section, we present a numerical study to verify and visualize our theoretical results. In order to compare MLE and SCE, we take $\bm{P}$ to be a reversible transition matrix. Besides, we adopt the following strategy to adjust the spectral gap of $\bm{P}$: for $10<\eta<$ 1 and a reversible transition matrix $\bm{P}$ with an invariant distribution $\mu$, we take
@@ Appendix A Comparison of different spectral gaps
This section provides a brief comparison of different spectral gaps introduced in Section 2. Since all of these results already appear in [5, 10], we present them here solely for the convenience of the reader. We first show that $\eta_{p}(\bm{P})\geq\eta_{s}(\bm{P})\geq\eta_{a}(\bm{P})$ and $\eta_{p}(\bm{P})\geq\eta_{ps}(\bm{P})/2$. As a direct consequence, our main results also holds if $\bm{P}$ admits a nonzero absolute spectral gap, a nonzero pseudo spectral gap,or a nonzero symmetric spectral gap.
@@
Chatterjee provided in [5] an example demonstrating that the ratio $\eta_{a}{\bm{P}}/\eta_{p}(\bm{P})$ can be arbitrarily small. Consider the one dimensional random walk on the periodic grid $\mathbb{Z}/n\mathbb{Z}$ where the probability of not moving and that of moving one step to the right are both $1/2$. It is then not hard to check that the IP gap of this Markov chain is of order $1/n$, while the symmetric spectral gap and the absolute spectral are both of order $1/n^{2}$.

### v3

@@ 3.3. Spectral properties of $\bm{P}_{2}$ and $\tilde{\bm{P}}_{2}$
$\mathbb{P}\left\{\text{choose }v\right\}=p(u,v),$ (3.24)
@@
$\frac{1}{4}\left(\bm{T}_{1}-\bm{S}_{1}^{\mathrm{T}}\right)\left(\bm{T}_{1}^{\mathrm{T}}-\bm{S}_{1}\right)=\frac{1}{2}\left(\bm{D}_{\mu}^{1/2}\frac{\bm{P}-\bm{\mathbf{I}}}{2}\bm{D}_{\mu}^{-1/2}+\bm{D}_{\mu}^{-1/2}\frac{\bm{P}^{\mathrm{T}}-\bm{\mathbf{I}}}{2}\bm{D}_{\mu}^{1/2}\right).$ (3.32)
@@
Similar to the proof for $\bm{P}_{2}$, we decompose $\tilde{\bm{P}}_{2}$ as the product of two transition matrices which represent step 1 and step 2 introduced in Subsection 3.1.2, respectively. To be specific, let $\tilde{\bm{S}}$ denote the transition matrix of step 1 in which a given state $\tilde{\bm{u}}\sim(u,v)$ will jump to $u$ or $v$ with the equal probability $1/2$, i.e. for any $\tilde{\bm{u}}\in\tilde{\Omega}^{2}$$u\in\Omega$,
@@
$\mathbb{P}\left\{\text{choose }v\right\}=p(u,v),$ (3.45)
@@ 4.1. Tail bounds of MLE and SCE
We will first derive tail bounds for the error of MLE and SCE using concentration inequalities for Markov chains. Let us state the complete version of [huang2024bernsteintypeinequalitiesmarkovchains, Theorem 2.4], which is a Bernstein-type inequality for Markov sums that only requires a nonzero IP gap of the transition matrix.
@@
In the above theorem, we use the symbol $\bm{A}$ for an arbitrary Markov transition kernel, so that it will not be confused with the particular transition matrix $\bm{P}$ in our context. Since $\eta_{p}(\bm{A})\geq\eta_{s}(\bm{A})\geq\eta_{a}(\bm{A})$ and $\eta_{p}(\bm{A})\geq\eta_{ps}(\bm{A})/2$, this concentration inequality also holds if the transition matrix admits a nonzero absolute spectral gap, a nonzero pseudo spectral gap,or a nonzero symmetric spectral gap. We then use Theorem 4.1 to derive non-asymptotic tail bounds for MLE.
@@
As for SCE, we choose to derive non-asymptotic tail bounds directly in the matrix form using matrix concentration inequalities. Recently, [neeman2024concentration] established Hoeffding and Bernstein-type matrix concentration inequalities (in $l_{2}$ operator norm) for sums of Markov dependent random matrices, under the assumption that the transition matrix admits a nonzero absolute spectral gap. We present a simplified version of [neeman2024concentration, Theorem 2.2] as follows.
@@ 4.2. Dimension-free expectation bound
Whether the error depends on the dimension shall rely on the choice of the error metric. The reason we can obtain dimension free expectation bound is that we measure the error using the Frobenius norm (i.e. $L_{2,u})$, which is a natural choice for matrices. If one instead uses the TV norm (i.e. $L_{1,u}$) to measure the error, a dependence on the dimension typically arises. This is why in [wolfer2021statistical], the error bound depends on the quantity $\|\bm{D}_{\mu}\bm{P}\|_{1/2}=(\sum_{u,v\in\Omega}(\mu(u)p(u,v))^{1/2})^{2}$, which may depend on the dimension $d$ and can be bounded by $d^{2}$.
@@
The estimate in (4.24) can be view as a matrix version of [chatterjee2025spectral, Theorem 1.2].In fact, one can apply [chatterjee2025spectral, Theorem 1.2] to $\{\bm{u}_{k}\}_{k\geq 1}$ to obtain that
@@
In this section, we present a numerical study to verify and visualize our theoretical results. In order to compare MLE and SCE, we take $\bm{P}$ to be a reversible transition matrix. Besides, we adopt the following strategy to adjust the spectral gap of $\bm{P}$: for $0<\eta<1$ and a reversible transition matrix $\bm{P}$ with an invariant distribution $\mu$, we take
@@ Appendix A Comparison of different spectral gaps
This section provides a brief comparison of different spectral gaps introduced in Section 2. Since all of these results already appear in [chatterjee2025spectral, huang2024bernsteintypeinequalitiesmarkovchains], we present them here solely for the convenience of the reader. We first show that $\eta_{p}(\bm{P})\geq\eta_{s}(\bm{P})\geq\eta_{a}(\bm{P})$ and $\eta_{p}(\bm{P})\geq\eta_{ps}(\bm{P})/2$. As a direct consequence, our main results also holds if $\bm{P}$ admits a nonzero absolute spectral gap, a nonzero pseudo spectral gap,or a nonzero symmetric spectral gap.
@@
Chatterjee provided in [chatterjee2025spectral] an example demonstrating that the ratio $\eta_{a}{\bm{P}}/\eta_{p}(\bm{P})$ can be arbitrarily small. Consider the one dimensional random walk on the periodic grid $\mathbb{Z}/n\mathbb{Z}$ where the probability of not moving and that of moving one step to the right are both $1/2$. It is then not hard to check that the IP gap of this Markov chain is of order $1/n$, while the symmetric spectral gap and the absolute spectral are both of order $1/n^{2}$.